A man has a daughter and a son.. The son is three years older than the daughter..
In one year the man will be six times as old as the daughter is now, and in ten years he will be fourteen years older than the combined ages of his children now..
What is the present age of the man..
PLease show me how you figured it out one tutor responded and said it was a negative number but i don't even know how to figure that out.

Answer by stanbon(75887)

NOW Data:
daughter = x yrs old.
son = x+3 yrs old
Dad = 6x-1
"IN TEN YEARS" Data:
Dad = (6x-1)+10= 6x+9 yrs old
EQUATION:
Dad in ten years = daughter now + son now + 14
6x+9=x+x+3+14
4x+9=17
4x=8
x=2
Dad's age now is 6x-1=11 yrs old
Highly unlikely.
Cheers,
Stan H.